# PGA: Personalizing Grasping Agents with Single Human-Robot Interaction

### v1

@@ III-D Personalized Object Grounding Model
The Personalized Object Grounding Model takes an image and a natural language indicator to infer the location of the personal object. We adopt the architecture of OFA [37], a Transformer [38] based Vision-Language model. The model’s optimization occurs by minimizing
$\mathcal{L}=-\sum_{\mathbf{o}^{\mathcal{K}}\in O^{\mathcal{K}}}logP_{\theta}(b_{\mathbf{o}^{\mathcal{K}}}|\mathcal{I}_{\mathbf{o}^{\mathcal{K}}},u_{\mathbf{o}^{\mathcal{K}}}),$ (3)
@@ III-E Personalized Object Grasping
IV EXPERIMENTS
IV-A Dataset
Our proposed dataset, GraspMine, is curated with 96 personal objects and 100+ everyday objects, encompassing Reminiscence images =$\mathcal{R}=\{I_{n}^{\mathcal{R}}\}_{n=1}^{N}$ and five distinct Test splits.
Reminiscence $\mathcal{R}$ involves $N=400$ raw images, each containing multiple objects. We annotated objects in the Reminiscence for analyzing the propagation ability of models and for training the Supervised method in Sec. IV-C and did not use the annotation in any process in training.
Test set is categorized into five distinct splits as follows. Each data comprises an image, a personal indicator, and its corresponding object bounding box (bbox) annotation.
  • Paraphrased split comprises all the images of Heterogeneous, Homogeneous, and Cluttered split with each personal indicator paraphrased by the annotators.
IV-B Compared Methods
GVCCI [2] automatically generates generic expressions of the objects in the Reminiscence and harnesses generated data for training. It shares the same model architecture with OFA.
Direct model is trained directly with the nodes $\mathbf{o}$ from the human-robot interaction in Sec. III-B. This model does not employ robot-object interaction in Sec. III-B and the Propagation through Reminiscence (Sec. III-C), resulting in training the model with a single sample per personal object.
Passive model is a variant of PGA, excluding robot-object interaction in Sec. III-B. In other words, Propagation through Reminiscence (Sec. III-C) is conducted only with the nodes from human-robot interaction in Sec. III-B.
Supervised model is trained with a fully-supervised approach that utilizes the ground-truth annotations (i.e., object location coordinates and personal indicators) in Reminiscence. It is regarded as the model showing the upper bound performance in the following experiments.
TABLE I: Results of offline experiments. Results show the personal object grounding score in GraspMine dataset. Bold scores represent the highest performance, while underlined scores the second-best results. The ‘interaction’ column represents the average number of interactions per object that users had with the robot to teach personal objects. ‘annotated’ indicates the total human annotations used in each method, while ‘utilized’ specifies the number of triplets - image, object bounding box, and personal indicator - employed for training the models. It’s noteworthy that subtracting the ‘annotated’ count from the ‘utilized’ count reveals the amount of triplet generated by each method.
@@ IV-B Compared Methods
IV-C Offline Experiment
@@
Evaluation Protocol. In the offline experiment, we study the PGA’s proficiency in grounding the target object given a personal indicator. By following standard practice [2, 37], we assess the Personalized Object Grounding Model using the Intersection over Union (IoU) score, a key visual grounding metric, which calculates the overlap between the predicted and ground truth bounding boxes. While we present the percentage of predictions surpassing an IoU of 0.5, we also emphasize IoU above 0.8 to ensure a stricter alignment between the predicted and actual regions, acknowledging the precision needed for successful object grasping. For the Cluttered split, we only reported scores with an IoU exceeding 0.8 since boxes with an IoU above 0.5 frequently included multiple objects, rendering the score less meaningful.
@@
Results. The offline results, presented in Tab. I, demonstrate that PGA substantially outperforms methods reliant on generic object knowledge, i.e., OFA and GVCCI. PGA’s advantage is also evident when compared to the Direct model, which solely uses object information acquired from human-robot interaction and ignores data from its Reminiscence. The improvement over the Passive model underscores the impact of robot-object interaction, emphasizing the importance of capturing objects from different views. Remarkably, even when compared to the Supervised method, which requires extensive annotations (91.3 annotations of per object on average) of Reminiscence, PGA, which only requires 1.1 interaction per object on average, shows comparable results. We also assess PGA in grounding objects when queried with generic instructions, using the Generic split. Remarkably, even after acquiring personal knowledge, PGA’s ability to ground objects through generic instructions remains impeccably intact. In short, by utilizing Reminiscence and the robot’s manipulative power, PGA proficiently grounds personal objects with learning from a single human-robot interaction, outperforming the baseline methods and showing comparable performance to the Supervised approach, all while retaining its knowledge of generic expressions.
@@
Impact of Reminiscence size. Shown in Fig. 3, our exploration delves into the relationship between PGA’s performance and the size of the Reminiscence — the number of raw images utilized. A size of ‘0’ is equivalent to the Direct model. We experimented on a log scale, examining sizes of 25, 100, and 400, with each configuration undergoing three separate trials. Our findings suggest a clear pattern: as PGA is exposed to an increasing number of scenes from the user’s environment, its performance is incrementally enhanced. This indicates that, as the robot collects more raw images from the environment, its performance is expected to be improved without necessitating further human supervision.
Analysis of Propagation through Noisy Reminiscence. Detecteed objects in the Reminiscence are noisy as they are predicted using an off-the-shelf object detector [20]. Therefore, in this analysis, we investigate how many noisy objects are pseudo-labelled with Propagation through Reminiscence. We divided noisy objects into two categories: 1) ambiguous boxes, i.e., boxes that only include a part of an object (e.g., Fig. 4-(a)) or boxes that surround multiple objects, and 2) invalid boxes, i.e., non-objects such as line stickers on desks or a robot arm (e.g., Fig. 4-B). We found 393 ambiguous boxes and 870 invalid boxes in the Reminiscence. We compare PGA with the Passive model that does not have robot-object interaction. Regarding propagation to ambiguous boxes, PGA showed a reduced rate of 41.2% compared with that of 50.4% from the Passive model. Moreover, Passive model had a high tendency to propagate towards invalid boxes, in at 83.7%, while PGA showed nearly negligible rate at 0.07%.
TABLE II: Results of Online Experiment (LCRG) in GraspMine.
@@
Supervised 100.0 83.3 83.3 70.0
Figure 4: Qualitative Analysis. The top row showcases examples of objects from each phase. In the bottom row, personalized object grounding results by PGA are depicted in blue boxes and those by Passive model in red, set alongside the ground truth in black boxes. Within the Propagation through Reminiscence on the top row, solid lined boxes are the pseudo-labelled objects from the Reminiscence and dotted boxes denote objects that were NOT pseudo-labelled with according indicators in the respective models.
IV-D Online Experiment
Results. The results, presented in Tab. II, reveal that PGA significantly surpassed the Direct model in grasping, along with enhanced grounding accuracy. The Direct model, which solely relies on object information acquired from human-robot interaction without leveraging raw images from Reminiscence, lagged behind PGA with a performance gap of 23.4 percent points in the Heterogeneous split and 33.3 percent points in the Homogeneous split. Notably, PGA’s performance was comparable to the Supervised method, which necessitates extensive labeling of raw Reminiscence images, highlighting our method’s efficacy with only requiring single human-robot interaction to transfer personal knowledge.
IV-E Qualitative Analysis
@@
Through qualitative analysis, shown in Fig. 4, we visualize examples from the various phases including Object Information Acquisition, Propagation through Reminiscence, and Personalized Object Grounding. For a comprehensive understanding, we compare the performance of PGA against the Passive model, which shows the most strong performance among the baselines. The Passive model does not leverage the data acquired from robot-object interaction.
@@
One striking observation is PGA’s adeptness at inspecting objects from diverse angles as seen in ‘robot-object interaction’. This allows PGA to recognize variations in object orientation, e.g., views from the rear or the base. In instances such as those represented in (c), (d), (e), and (f), the Passive model often overlooks objects in the Reminiscence that differ from their presentations in the human-robot interaction phase. We conjecture that the omitted samples, e.g., (d), (e), and (f), in the training phase affect some wrong prediction results as in C and D. Moreover, as detailed in Sec. IV-C, the Passive model occasionally tends to propagate towards ambiguous boxes in the Reminiscence, as depicted in (a). Such behavior can lead to inaccurate grounding, as observed in A. The tendency of propagating towards incorrect boxes might have affected in some results shown in panel B.
@@
Despite PGA’s robust grounding abilities as seen through A-D, challenges persist in some complex scenarios, such as in a scenario like E, where objects are situated in highly cluttered environments. Nevertheless, our qualitative analysis reveals PGA’s successful Propagation through Reminiscence, enhanced by the robot-object interaction, yielding promising results in Personalized Object Grounding.
V CONCLUSION

### v2

@@ III-D Personalized Object Grounding Model
The Personalized Object Grounding Model is a Transformer [36] based Vision-Language model that takes an image and a natural language indicator to infer the bounding box coordinates of the personal object. The model’s optimization occurs by minimizing
$\mathcal{L}=-\sum_{\mathbf{o}^{\mathcal{K}}\in O^{\mathcal{K}}}logP_{\theta}(b_{\mathbf{o}^{\mathcal{K}}}|I_{\mathbf{o}^{\mathcal{K}}},u_{\mathbf{o}^{\mathcal{K}}}),$ (3)
@@ III-E Personalized Object Grasping
TABLE I: Results of offline experiments. Results show the personal object grounding score in GraspMine dataset. Bold scores represent the highest performance, while underlined scores the second-best results. The ‘interaction’ column represents the average number of interactions per object that users had with the robot to teach personal objects. ‘annotated’ indicates the total human annotations used in each method, while ‘utilized’ specifies the number of triplets - image, object bounding box, and personal indicator - employed for training the models. It’s noteworthy that subtracting the ‘annotated’ count from the ‘utilized’ count reveals the amount of triplet automatically generated by each method without any human labor.
@@
IV GraspMine
Our proposed task scenario, GraspMine, consists of a curated dataset comprising a Training set, Reminiscence, and Test set, featuring 96 personal objects along with 100+ everyday objects, totaling around 200 individual objects.
Training set consists of 96 pairs of images $I_{o}$ and their respective personal indicators $u_{o}$, alongside general indicators $u_{o}^{G}$, collected through human-robot interaction described in Sec. III-B and illustrated in Fig. 2-(b).
Reminiscence, =$\mathcal{R}=\{I_{n}^{\mathcal{R}}\}_{n=1}^{N}$, involves $N=400$ raw images, each containing multiple objects. These unlabeled images can be optionally provided to aid the learning process. However, we annotated objects in the Reminiscence for analyzing the propagation ability of models and for training the Supervised method in Sec. V-B. Notably, these annotations were not utilized in any aspect of training PGA.
[…]
  • Paraphrased split comprises all Heterogeneous, Homogeneous, and Cluttered split with each personal indicator paraphrased by the annotators. For instance, the personal indicator “my sleeping pills” is queried as “the medication I take for sleep” as depicted in Fig. 1.
V EXPERIMENTS
V-A Compared Methods
GVCCI [2] is a robotic lifelong learning framework designed to autonomously learn generic expressions of objects. By leveraging images from both the training set and the Reminiscence, it automatically generates these generic expressions and utilizes the generated data for training.
Direct model is trained directly from the human-robot interaction (Sec. III-B). This model does not employ robot-object interaction (Sec. III-B) and the Propagation through Reminiscence (Sec. III-C), resulting in training the model with a single sample per personal object. This model serves as a representation of traditional LCRG systems that typically learn in a supervised manner when confronted with the task scenario of GraspMine. Consequently, the Direct model can only utilize 96 annotated data obtained from an average of 1.1 human-robot interaction per object (101 interactions for 96 objects due to five visual grounding failures).
PassivePGA model is an ablative model of PGA, excluding robot-object interaction in Sec. III-B. In other words, Propagation through Reminiscence (Sec. III-C) is conducted only with the nodes from human-robot interaction in Sec. III-B. Consequently, this model utilizes 96 annotated data points along with 4.7k pseudo-labeled data obtained from the Propagation through Reminiscence, resulting in a total of 4.8k utilized data points.
Supervised model is trained with a fully-supervised approach that utilizes ground-truth annotations (i.e., object location coordinates and personal indicators) in Reminiscence. Since this model has access to all 8.7k ground truth annotations in the Reminiscence, whereas GraspMine assumes access of only one annotation per object (96 annotations in total), the Supervised model serves as the upper bound performance benchmark for GraspMine in the following experiments. Essentially, training this model requires approximately 100 times more manual effort compared to PGA.
V-B Offline Experiment
@@
Evaluation Protocol. In the offline experiment, we study the PGA’s proficiency in grounding the target object given a personal indicator. By following standard practice [2, 39], we assess the grounding accuracy using the Intersection over Union (IoU) score, a key visual grounding metric, which calculates the overlap between the predicted and ground truth bounding boxes. While we present the percentage of predictions surpassing an IoU of 0.5, we also emphasize IoU above 0.8 to ensure a stricter alignment between the predicted and actual regions, acknowledging the precision needed for successful object grasping. For the Cluttered split, we only reported scores with an IoU exceeding 0.8 since boxes with an IoU above 0.5 frequently included multiple objects, rendering the score less meaningful.
@@
Comparison with baselines. The offline results, presented in Tab. I, demonstrate that the Direct model, which relies solely on object information acquired from human-robot interaction without leveraging Reminiscence, exhibited only marginal improvement compared to methods reliant on generic object knowledge, i.e., OFA and GVCCI. This indicates that naively training existing LCRG models on a small amount of labeled data does not make significant performance gains in GraspMine. However, PGA, utilizing a large amount of unlabeled data, demonstrated a significant improvement compared to the Direct model, achieving approximately 30% enhancement. Remarkably, even when compared to the Supervised method, which requires about 100 times more annotations than PGA, PGA exhibited comparable results. We also assess PGA in grounding objects when queried with generic instructions, using the Generic split, to examine how well PGA retains knowledge about the generic instructions after training the personal indicators. Even after acquiring personal knowledge, PGA’s ability to ground objects through generic instructions remains impeccably intact. In short, by utilizing an Reminiscence and the robot’s manipulative power, PGA proficiently grounds personal objects with learning from a single human-robot interaction, outperforming the baseline methods and showing comparable performance to the Supervised approach, all while retaining its knowledge of generic expressions.
Impact of robot-object interaction. Tab. I shows a notable enhancement of up to 8.5% in PGA over the PassivePGA model. This underscores the impact of robot-object interaction in improving performance by capturing objects from different views. Notably, when comparing the $IoU_{>0.5}$ and $IoU_{>0.8}$ across various splits in Tab. I, we found that the improvement in $IoU_{>0.8}$ (+5-8%) exceeded that of $IoU_{>0.5}$ (+1-2%). This suggests that leveraging robot-object interaction leads to more precise localization of objects, particularly beneficial for robotic grasping. To further understand these findings, we compare the propagation accuracy of PassivePGA and PGA. The accuracy, measured as the percentage of correctly pseudo-labeled objects among total nodes, revealed PGA achieving 79.4%, while PassivePGA lagged behind at 64.5%. We further investigate incorrectly labeled objects and observe that most of the propagation error comes from the noisiness of the off-the-shelf object detector [19]. Noisy objects were categorized into two groups: 1) ambiguous boxes, i.e., boxes that only include a part of an object (e.g., Fig. 4-(a)) or boxes that surround multiple objects, and 2) invalid boxes, i.e., non-objects such as line stickers on desks or a robot arm (e.g., Fig. 4-B). We found 393 ambiguous boxes and 870 invalid boxes in the Reminiscence. Regarding propagation to ambiguous boxes, PGA showed a reduced rate of 41.2% compared with that of 50.4% from the PassivePGA model. Moreover, PassivePGA model had a high tendency to propagate towards invalid boxes, at 83.7%, while PGA showed a nearly negligible rate at 0.07%. In summary, the impact of robot-object interaction on the inaccurate and noisy Propagation through Reminiscence appears to have influenced both the overall grounding accuracy and the precision of localization.
@@
Impact of Reminiscence size. As shown in Fig. 3, our exploration delves into the relationship between PGA’s performance and the size of the Reminiscence — the number of raw images utilized. A size of ‘0’ is equivalent to the Direct model. We experimented on a log scale, examining sizes of 25, 100, and 400, with each configuration undergoing three separate trials. Our findings suggest a clear pattern: as PGA is exposed to an increasing number of scenes from the user’s environment, its performance is incrementally enhanced. This indicates that, as the robot collects more raw images from the environment, its performance is expected to be improved without necessitating further human supervision.
Figure 4: Qualitative Analysis. The top row showcases examples of objects from each phase. In the bottom row, personalized object grounding results by PGA are depicted in blue boxes and those by PassivePGA model in red, set alongside the ground truth in black boxes. Within the Propagation through Reminiscence on the top row, solid lined boxes are the pseudo-labeled objects from the Reminiscence and dotted boxes denote objects that were NOT pseudo-labeled according to indicators in the respective models.
V-C Online Experiment
Results. Results in Tab. II reveal that PGA significantly surpassed the Direct model in grasping, along with enhanced grounding accuracy. The Direct model, which solely relies on object information acquired from human-robot interaction without leveraging raw images from Reminiscence, lagged behind PGA with a performance gap of 23.4% in the Heterogeneous split and 33.3% in the Homogeneous split. This result underscores the effectiveness of PGA, which achieves superior performance without requiring additional human labor for annotations compared to the Direct model.
TABLE II: Results of Real-world Online Experiment (LCRG) in GraspMine.
@@ V-C Online Experiment
V-D Qualitative Analysis
@@
Through qualitative analysis, shown in Fig. 4, we visualize examples from the various phases including Object Information Acquisition, Propagation through Reminiscence, and Personalized Object Grounding. For a comprehensive understanding, we compare the performance of PGA against the PassivePGA model.
@@
One striking observation is PGA’s adeptness at inspecting objects from diverse angles as seen in ‘robot-object interaction’. This allows PGA to recognize variations in object orientation, e.g., views from the rear or the base. In instances such as those represented in (c), (d), (e), and (f), the PassivePGA model often overlooks objects in the Reminiscence that differ from their appearances in the human-robot interaction phase. We conjecture that the omitted samples, e.g., (d), (e), and (f), in the training phase affect some wrong prediction results as in C and D. For example, consider the case of “the charger for my brand new MacBook”. With the aid of robot-object interaction, PGA successfully propagates to samples viewed from different angles. However, PassivePGA fails to capture such samples as shown in (d) and (e). In sample D, it is evident that PassivePGA makes an incorrect inference, while PGA’s inference is accurate. Moreover, as detailed in Sec. V-B, the PassivePGA model occasionally tends to propagate towards ambiguous boxes in the Reminiscence, as depicted in (a). Such behavior can lead to inaccurate grounding, as observed in the red box in sample A. The tendency of propagating towards incorrect boxes might have affected in some results shown in panel B.
@@
Despite PGA’s robust grounding abilities as seen through A-D, challenges persist in some complex scenarios, such as in a scenario like E, where objects are situated in highly cluttered environments—the difficulty reflected in the lower scores of Cluttered split in Tab. I. Nevertheless, our qualitative analysis reveals PGA’s successful Propagation through Reminiscence, enhanced by the robot-object interaction, yielding promising results in Personalized Object Grounding.
VI DISCUSSION
The Personalized Grasping Agent (PGA) has shown promising performance in the GraspMine dataset. However, it is essential to discuss several assumptions and limitations in our experimental setup to guide future studies for better improvement in dealing with GraspMine.
Due to the use of an arm robot without mobility necessitates two primary assumptions. Firstly, we confine the execution area to a 80*60 cm table where all objects are at least partially visible within this space. This assumption limits manipulation to flat and open surfaces, excluding scenarios such as retrieving objects from handbags or shelves. Overcoming this limitation would require developing deeper and dynamically adaptive reasoning in motion planning, a direction we leave for future exploration. Secondly, we assume a full access to raw images of Reminiscence. While obtaining these raw images from a mobile robot through random navigation in the user’s environment is feasible, we assumed that these images were pre-acquired. Extending this work with a mobile robot capable of autonomously acquiring raw images from the user’s environment would be a crucial topic for further studies.
During the online experiment with the physical robot (See Tab. II), a performance drop from grounding accuracy to grasping accuracy (around 13%) was observed. This decline can be attributed to limitations in motion planning, particularly in the manipulation strategy employed, which first aligns the object and the end effector’s horizontal coordinates before making vertical adjustments. However, this approach may prove ineffective when attempting to grasp objects like wine bottles. Enhancing motion planning to tailor grasping strategies for diverse objects is crucial for future studies.
VII CONCLUSION